OpenStudy (anonymous):

5.) which procedure could be used to graph the line represented by the following equation? Y=1/2x+3 (1 point) A. Plot x intercept (3,0) and use slope of 1/2 to find a second point B. plot y intercept (0,3) and use the slope of 1/2 to find a second point C. Plot x intercept (1/2,0) and use slope of 3 to find a second point D. Plot y intercept (0,1/2) and use slope of 3 to find a second point 6.) which of the following equations best represents the regression line for the data given in the table above? X:(1,2,4,6,8,10,11) Y: (0,-1,0,4,8,9,8) A.y=x+2 B: y=2x-2 C.y=-x-2 D.y=x-2 7.)Which eq

OpenStudy (texaschic101):

Let me explain this a little bit... To find the slope and the y intercept...put your equation in y = mx + b form, which in this problem, already is. The number in the m position is the slope. So the slope is 1/2. The number in the b position is the y intercept. So the y intercept is (0,3). The x intercept can be found by subbing in 0 for y and solving for x.. y = 1/2x + 3 0 = 1/2x + 3 -3 = 1/2x -3/(1/2) = x -3 * 2 = x -6 = x so the x intercept is (-6,0) so our results for y = 1/2x + 3 are : slope = 1/2 y intercept is (0,3) x intercept is (-6,0) so that eliminates all but 1 of the answers and I think you can figure it out. ======================